Find the mean of the probability distribution of a random variable X, if P(X) = 20

X 1

for X =

1,2,3,4, and 5.


1
Expert's answer
2022-03-15T05:42:49-0400

Given that P(X) = "\frac{X+1}{20}"

"E(X)=\sum(X.P(X))"

= 1 p(1) + 2 p(2) + 3 p(3) + 4 p(4) + 5 p(5)


= (1)(2/20) + (2)(3/20) + (3)(4/20) + (4)(5/20) + (5)(6/20)

=2/20 + 6/20 + 12/20 + 20/20 + 30/20 =70/2

= 3.5.

Answer: 3.5